OpenStudy (disco619):

How do I solve this? Help? 2/3 of a plot of Land is garden. 1/5 of the garden is lawn. Find the fraction of the Plot of Land which is lawn.

OpenStudy (bibby):

Get them into like terms Use 15 as a denominator. Then the plot of land is divided into 15 parts

OpenStudy (disco619):

U are close... I have the Answer but don't know how it came

OpenStudy (disco619):

Answer is \[ \frac{ 2}{ 15 }\] But HOW?

OpenStudy (disco619):

LOL Nevermind. I guess we had to Multiply these Both.

OpenStudy (disco619):

But still don't Understand why....

OpenStudy (bibby):

oh wait I misunderstood So 2/3 of the land is garden 1/5 of that is lawn Yeah multiply

OpenStudy (disco619):

But why? What's the Reason?

OpenStudy (bibby):

You're taking a fraction OF a fraction. Pretend you have a box. You cut it in 2. Now you cut one half in 2. You're left with 1/4

OpenStudy (disco619):

Huh, didn't Understood but Good enough for me. I have no hope in Maths. Tomorrow is my Exam xD Thanks... P.S: If it is a Fraction of a Fraction, then wouldn't we Divide these Fractions with each other rather Multiply them?

OpenStudy (bibby):

lol, trying to think of an actual way to explain this as it's just something I "know" I'll try and help you visualize it

OpenStudy (disco619):

Yea, i was about to Make that SAME drawing. I think i get it, if an Equation matching this kind of Situation, i will know to Multiply them with each other. Thanks again!

OpenStudy (bibby):

Eh? Equations? Gimme an example question as I don't think I get that

OpenStudy (disco619):

I mean, for example the Question is... Q] 2/5 of a pizza has Pepper. 4/5 of the pizza with pepper has salt. Find the fraction of the pizza with Salt.

OpenStudy (bibby):

just wondering, are you typing something or should I answer that question

OpenStudy (disco619):

then we will do this? \[\frac{ 2 }{ 5 } \times \frac{ 4 }{ 5 } \] \[\frac{ 8 }{ 10 } \] \[\frac{ 4 }{ 5 } \] ???

OpenStudy (bibby):

you seem to have made an error, damn thing just glitched out...